# On degenerate metric resolvent: well-definedness and new convergence analysis

### v3

@@ 2.2 Basic properties
(i) $\mathcal{Q}=\mathcal{Q}\circ\mathcal{P}_{\mathsf{ran}\mathcal{Q}}=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{Q}$;
(ii) $\mathcal{T}\circ\mathcal{P}_{\mathsf{ran}\mathcal{Q}}=\mathcal{T}$, i.e., $\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}=\mathcal{T}$.
Now, we show several basic properties on $\mathcal{T}$ based on Assumption 2 and Fact 1. The first is the single-valuedness of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$.
@@
Under Assumptions 1 and 2, $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$ is single-valued.
Proof 2.4.
Let $x\in\mathsf{dom}\mathcal{T}$ and assume that $y_{1}\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$ and $y_{2}\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$. Then, we have
@@
which yields $\|y_{1}-y_{2}\|_{\mathcal{Q}}=0$, i.e., $\mathcal{Q}y_{1}=\mathcal{Q}y_{2}$, or equivalently, $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}y_{1}=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}y_{2}$.
[7, Proposition 4.2] and [11, Lemma 3.2] showed that $\mathcal{T}$ is $\mathcal{Q}$-firmly nonexpansive under maximally monotone $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ and non-degenerate $\mathcal{Q}$. This was also discussed in [32, Sect. 2]. As emphasized in [10, Lemma 2.6], this key property is preserved under restricted monotonicity of $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ for degenerate case.
@@
[10, Lemma 2.6] Under Assumptions 1 and 2, $\mathcal{T}$ given as (2) is $\mathcal{Q}$-firmly nonexpansive, i.e.,
$\big{\|}\mathcal{T}x-\mathcal{T}y\big{\|}_{\mathcal{Q}}^{2}\leq\big{\langle}\mathcal{T}x-\mathcal{T}y\big{|}\mathcal{Q}(x-y)\big{\rangle},\quad\forall x,y\in\mathsf{dom}\mathcal{T},$
@@
$\big{\|}\mathcal{T}x-\mathcal{T}y\big{\|}_{\mathcal{Q}}^{2}+\big{\|}(\mathcal{I}-\mathcal{T})x-(\mathcal{I}-\mathcal{T})y\big{\|}_{\mathcal{Q}}^{2}\leq\big{\|}x-y\big{\|}_{\mathcal{Q}}^{2},\quad\forall x,y\in\mathsf{dom}\mathcal{T}.$ (6)
@@
There is still a slight difference between our result and [10, Lemma 2.6]. The latter required $\mathcal{T}$ to be well-defined, which, however, is not assumed in our result. Nevertheless, we clarify that all the expressions in above inequalities are well-defined, even if $\mathcal{T}$ is multi-valued. Indeed, $\|\mathcal{T}x\|_{\mathcal{Q}}=\|\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{T}x\|_{\mathcal{Q}}$ and $\langle\mathcal{T}x|\mathcal{Q}y\rangle=\langle\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{T}x|\mathcal{Q}y\rangle$, $\forall(x,y)\in\mathsf{dom}\mathcal{T}\times\mathcal{H}$. They are well-defined, since $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$ is single-valued (by Lemma 2.3). To put it more explicitly, combining with Fact 1, one can equivalently rewrite (6) as $\forall x_{\mathsf{r}},y_{\mathsf{r}}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{dom}\mathcal{T})$:
$\big{\|}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}})-(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(y_{\mathsf{r}})\big{\|}_{\mathcal{Q}}^{2}+\big{\|}(\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}})-(\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(y_{\mathsf{r}})\big{\|}_{\mathcal{Q}}^{2}\leq\big{\|}x_{\mathsf{r}}-y_{\mathsf{r}}\big{\|}_{\mathcal{Q}}^{2}.$ (7)
The concept of demiclosedness [4, Definition 4.26] and Browder’s demiclosedness principle [4, Theorem 4.27] are crucial for convergence analysis, see [11, Lemma 3.4, Theorem 3.1] for example. Here it is natural (or evenly, trivial) to extend the results to any $\mathcal{Q}$-firmly nonexpansive operator $\mathcal{S}$.
@@
Under Assumption 1, let $D$ be a nonempty weakly sequentially closed subset of $\mathsf{ran}\mathcal{Q}$, let $\mathcal{S}:D\mapsto\mathsf{ran}\mathcal{Q}$, and let $u\in\mathsf{ran}\mathcal{Q}$. Then $\mathcal{S}$ is $\mathcal{Q}$-demiclosed at $u$ if, for every sequence $\big{\{}x^{k}\big{\}}_{k\in\mathbb{N}}$ in $D$ and every $x\in D$ such that $x^{k}\rightharpoonup x$ and $\mathcal{S}x^{k}\rightarrow u$, we have $\mathcal{S}x=u$. In addition, $\mathcal{S}$ is $\mathcal{Q}$-demiclosed if it is demiclosed at every point in $D$.
@@
Under Assumption 1, let $D$ be a nonempty weakly sequentially closed subset of $\mathsf{ran}\mathcal{Q}$ and let $\mathcal{S}:D\mapsto\mathsf{ran}\mathcal{Q}$ be a $\mathcal{Q}$-nonexpansive operator, i.e.,
$\big{\|}\mathcal{S}x_{1}-\mathcal{S}x_{2}\big{\|}_{\mathcal{Q}}\leq\big{\|}x_{1}-x_{2}\big{\|}_{\mathcal{Q}},\quad\forall x_{1},x_{2}\in D\subseteq\mathsf{ran}\mathcal{Q}\cap\mathsf{dom}\mathcal{S}.$
@@
Lemma 2.8 ($\mathcal{Q}$-demiclosedness of $\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}$).
Given $\mathcal{T}$ as (2) satisfying Assumptions 1 and 2, let $D$ be a nonempty closed subset of $\mathsf{ran}\mathcal{Q}\cap\mathsf{dom}\mathcal{T}$, then $\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}:D\mapsto\mathsf{ran}\mathcal{Q}$ is $\mathcal{Q}$-demiclosed.
Proof 2.9.
In view of Lemma 2.7, it suffices to show that $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}:D\mapsto\mathsf{ran}\mathcal{Q}$ is $\mathcal{Q}$-nonexpansive. Indeed, this immediately follows from Fact 2.
@@ 2.3 Generalized Moreau’s decomposition identity
Lemma 2.10 develops some equivalent forms of $\mathcal{T}$, which generalize the classic Moreau’s decomposition identity [4, Propositioin 23.20], and will be used for proving Theorem 3.1. The proof is postponed to Appendix A.
@@
where $\sqrt{\mathcal{Q}}$ is defined as Fact 1-(iii).
@@ 3.1 Main result
Under Assumption 1, $\mathcal{T}$ given as (2) has full domain, if and only if $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ (or equivalently saying, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}$) $:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$ is maximally monotone.
Let us first discuss the maximal monotonicity of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$, from which then readily follows the proof of Theorem 3.1. This concept is essentially a restricted version of maximal monotonicity from $\mathcal{H}$ to $2^{\mathcal{H}}$, by confining both domain and range of the operator to $\mathsf{ran}\mathcal{Q}$. Indeed, this restricted maximal monotonicity is closely related to Definition 2.1, due to the following lemma.
@@
The following 3 statements are equivalent.
(i) $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ is monotone;
(ii) $\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone;
(iii) $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone.
@@
(ii)$\Longrightarrow$(iii): If $\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone, then we have:
$\langle u-v|x-y\rangle\geq 0,\ \forall(u,x)\in\mathsf{gra}(\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}),\ \forall(v,y)\in\mathsf{gra}(\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}).$
Note that $\langle u-v|x-y\rangle=\langle u-v|x_{\mathsf{r}}-y_{\mathsf{r}}\rangle$, since $u,v\in\mathsf{ran}\mathcal{Q}$. This is equivalent to saying that
$\langle u-v|x_{\mathsf{r}}-y_{\mathsf{r}}\rangle\geq 0,\ \forall(u,x_{\mathsf{r}})\in\mathsf{gra}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}),\ \forall(v,y_{\mathsf{r}})\in\mathsf{gra}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}).$
This shows that $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone.
(iii)$\Longrightarrow$(ii): If $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone, then we have:
$\langle u-v|\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}y\rangle\geq 0,\ \forall(u,x)\in\mathsf{gra}(\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}),\ \forall(v,y)\in\mathsf{gra}(\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}).$
Note that $\langle u-v|\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}y\rangle=\langle\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(u-v)|x-y\rangle=\langle u-v|x-y\rangle\geq 0$, since $u,v\in\mathsf{ran}\mathcal{Q}$. This shows that $\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone.
It is natural to develop a restricted version of classic Minty’s theorem [4, Theorem 21.1] for the restricted maximality of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$.
@@
Let $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ be monotone. Then $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone, if and only if $\mathsf{ran}(\mathcal{I}+\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}})=\mathsf{ran}\mathcal{Q}$.
The proof is exactly the same as [4, Theorem 21.1], if the subspace $\mathsf{ran}\mathcal{Q}$ is regarded as the ambient space. Actually, our specific $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ in Theorem 3.4 can be replaced by any monotone operator $\mathcal{B}:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$, of course, including $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}$.

### v4

@@ 2.2 Basic properties
(i) $\mathcal{Q}=\mathcal{Q}\circ\mathcal{P}_{\mathrm{ran}\mathcal{Q}}=\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{Q}$;
(ii) $\mathcal{T}\circ\mathcal{P}_{\mathrm{ran}\mathcal{Q}}=\mathcal{T}$, i.e., $\mathcal{T}|_{\mathrm{ran}\mathcal{Q}}=\mathcal{T}$.
Now, we show several basic properties on $\mathcal{T}$ based on Assumption 2 and Fact 1. The first is the single-valuedness of $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T}$.
@@
Under Assumptions 1 and 2, $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T}$ is single-valued.
Proof 2.5.
Let $x\in\mathrm{dom}\mathcal{T}$ and assume that $y_{1}\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$ and $y_{2}\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$. Then, we have
@@
which yields $\|y_{1}-y_{2}\|_{\mathcal{Q}}=0$, i.e., $\mathcal{Q}y_{1}=\mathcal{Q}y_{2}$, or equivalently, $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}y_{1}=\mathcal{P}_{\mathrm{ran}\mathcal{Q}}y_{2}$.
[8, Proposition 4.2] and [15, Lemma 3.2] showed that $\mathcal{T}$ is $\mathcal{Q}$-firmly nonexpansive under maximally monotone $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ and non-degenerate $\mathcal{Q}$. This was also discussed in [50, Sect. 2]. As emphasized in [13, Lemma 2.6], this key property is preserved under restricted monotonicity of $\mathrm{gra}\mathcal{A}\cap(\mathcal{H}\times\mathrm{ran}\mathcal{Q})$ for degenerate case.
@@
[13, Lemma 2.6] Under Assumptions 1 and 2, $\mathcal{T}$ given as (2) is $\mathcal{Q}$-firmly nonexpansive, i.e.,
$\big{\|}\mathcal{T}x-\mathcal{T}y\big{\|}_{\mathcal{Q}}^{2}\leq\big{\langle}\mathcal{T}x-\mathcal{T}y\big{|}\mathcal{Q}(x-y)\big{\rangle},\quad\forall x,y\in\mathrm{dom}\mathcal{T},$
@@
$\big{\|}\mathcal{T}x-\mathcal{T}y\big{\|}_{\mathcal{Q}}^{2}+\big{\|}(\mathcal{I}-\mathcal{T})x-(\mathcal{I}-\mathcal{T})y\big{\|}_{\mathcal{Q}}^{2}\leq\big{\|}x-y\big{\|}_{\mathcal{Q}}^{2},\quad\forall x,y\in\mathrm{dom}\mathcal{T}.$ (6)
@@
There is still a slight difference between our result and [13, Lemma 2.6]. The latter required $\mathcal{T}$ to be well-defined, which, however, is not assumed in our result. Nevertheless, we clarify that all the expressions in above inequalities are well-defined, even if $\mathcal{T}$ is multi-valued. Indeed, $\|\mathcal{T}x\|_{\mathcal{Q}}=\|\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{T}x\|_{\mathcal{Q}}$ and $\langle\mathcal{T}x|\mathcal{Q}y\rangle=\langle\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{T}x|\mathcal{Q}y\rangle$, $\forall(x,y)\in\mathrm{dom}\mathcal{T}\times\mathcal{H}$. They are well-defined, since $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T}$ is single-valued (by Lemma 2.4). To put it more explicitly, combining with Fact 1, one can equivalently rewrite (6) as $\forall x_{\mathsf{r}},y_{\mathsf{r}}\in\mathcal{P}_{\mathrm{ran}\mathcal{Q}}(\mathrm{dom}\mathcal{T})$:
$\big{\|}(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}})-(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})(y_{\mathsf{r}})\big{\|}_{\mathcal{Q}}^{2}+\big{\|}(\mathcal{I}-\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}})-(\mathcal{I}-\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})(y_{\mathsf{r}})\big{\|}_{\mathcal{Q}}^{2}\leq\big{\|}x_{\mathsf{r}}-y_{\mathsf{r}}\big{\|}_{\mathcal{Q}}^{2}.$ (7)
The concept of demiclosedness [5, Definition 4.26] and Browder’s demiclosedness principle [5, Theorem 4.27] are crucial for convergence analysis, see [15, Lemma 3.4, Theorem 3.1] for example. Here it is natural (or evenly, trivial) to extend the results to any $\mathcal{Q}$-firmly nonexpansive operator $\mathcal{S}$.
@@
Under Assumption 1, let $D$ be a nonempty weakly sequentially closed subset of $\mathrm{ran}\mathcal{Q}$, let $\mathcal{S}:D\mapsto\mathrm{ran}\mathcal{Q}$, and let $u\in\mathrm{ran}\mathcal{Q}$. Then $\mathcal{S}$ is $\mathcal{Q}$-demiclosed at $u$ if, for every sequence $\big{\{}x^{k}\big{\}}_{k\in\mathbb{N}}$ in $D$ and every $x\in D$ such that $x^{k}\rightharpoonup x$ and $\mathcal{S}x^{k}\rightarrow u$, we have $\mathcal{S}x=u$. In addition, $\mathcal{S}$ is $\mathcal{Q}$-demiclosed if it is demiclosed at every point in $D$.
@@
Under Assumption 1, let $D$ be a nonempty weakly sequentially closed subset of $\mathrm{ran}\mathcal{Q}$ and let $\mathcal{S}:D\mapsto\mathrm{ran}\mathcal{Q}$ be a $\mathcal{Q}$-nonexpansive operator, i.e.,
$\big{\|}\mathcal{S}x_{1}-\mathcal{S}x_{2}\big{\|}_{\mathcal{Q}}\leq\big{\|}x_{1}-x_{2}\big{\|}_{\mathcal{Q}},\quad\forall x_{1},x_{2}\in D\subseteq\mathrm{ran}\mathcal{Q}\cap\mathrm{dom}\mathcal{S}.$
@@
Lemma 2.9 ($\mathcal{Q}$-demiclosedness of $\mathcal{I}-\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathrm{ran}\mathcal{Q}}$).
Given $\mathcal{T}$ as (2) satisfying Assumptions 1 and 2, let $D$ be a nonempty closed subset of $\mathrm{ran}\mathcal{Q}\cap\mathrm{dom}\mathcal{T}$, then $\mathcal{I}-\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathrm{ran}\mathcal{Q}}:D\mapsto\mathrm{ran}\mathcal{Q}$ is $\mathcal{Q}$-demiclosed.
Proof 2.10.
In view of Lemma 2.8, it suffices to show that $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathrm{ran}\mathcal{Q}}:D\mapsto\mathrm{ran}\mathcal{Q}$ is $\mathcal{Q}$-nonexpansive. Indeed, this immediately follows from Fact 2.
@@ 2.3 Generalized Moreau’s decomposition identity
Lemma 2.11 develops some equivalent expressions, which generalize the classic Moreau’s decomposition identity [5, Propositioin 23.20], and will be used for proving Theorem 3.1. The proof is postponed to Appendix A.
@@
where $\sqrt{\mathcal{Q}}$ is defined as Fact 1-(iii), the numberings (1)–(4) are used in the proof.
@@ 3.1 Main result
Under Assumptions 1 and 2, $\mathcal{T}$ given as (2) has full domain, if and only if $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}$ (or equivalently saying, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathrm{ran}\mathcal{Q}}$) $:\mathrm{ran}\mathcal{Q}\mapsto 2^{\mathrm{ran}\mathcal{Q}}$ is maximally monotone.
Before proving Theorem 3.1, let us first discuss the maximal monotonicity of $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}:\mathrm{ran}\mathcal{Q}\mapsto 2^{\mathrm{ran}\mathcal{Q}}$. This is essentially a restricted version of maximal monotonicity from $\mathcal{H}$ to $2^{\mathcal{H}}$, by confining both domain and range of the operator to $\mathrm{ran}\mathcal{Q}$. Indeed, this restricted maximal monotonicity is closely related to Definition 2.1, due to the following lemma.
@@
The following three statements are equivalent.
(i) $\mathrm{gra}\mathcal{A}\cap(\mathcal{H}\times\mathrm{ran}\mathcal{Q})$ is monotone;
(ii) $\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}$ is monotone;
(iii) $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}$ is monotone.
@@
(ii)$\Longrightarrow$(iii): If $\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}$ is monotone, then we have:
$\langle u-v|x-y\rangle\geq 0,\ \forall(u,x)\in\mathrm{gra}(\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}),\ \forall(v,y)\in\mathrm{gra}(\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}).$
Note that $\langle u-v|x-y\rangle=\langle u-v|x_{\mathsf{r}}-y_{\mathsf{r}}\rangle$, since $u,v\in\mathrm{ran}\mathcal{Q}$. This is equivalent to saying that
$\langle u-v|x_{\mathsf{r}}-y_{\mathsf{r}}\rangle\geq 0,\ \forall(u,x_{\mathsf{r}})\in\mathrm{gra}(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}),\ \forall(v,y_{\mathsf{r}})\in\mathrm{gra}(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}).$
This shows that $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}$ is monotone.
(iii)$\Longrightarrow$(ii): If $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}$ is monotone, then we have:
$\langle u-v|\mathcal{P}_{\mathrm{ran}\mathcal{Q}}x-\mathcal{P}_{\mathrm{ran}\mathcal{Q}}y\rangle\geq 0,\ \forall(u,x)\in\mathrm{gra}(\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}),\ \forall(v,y)\in\mathrm{gra}(\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}).$
Note that $\langle u-v|\mathcal{P}_{\mathrm{ran}\mathcal{Q}}x-\mathcal{P}_{\mathrm{ran}\mathcal{Q}}y\rangle=\langle\mathcal{P}_{\mathrm{ran}\mathcal{Q}}(u-v)|x-y\rangle=\langle u-v|x-y\rangle\geq 0$, since $u,v\in\mathrm{ran}\mathcal{Q}$. This shows that $\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}$ is monotone.
It is natural to develop a restricted version of classic Minty’s theorem [5, Theorem 21.1] for the restricted maximality of $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}$.
@@
Let $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}$ be monotone. Then $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}$ is maximally monotone, if and only if $\mathrm{ran}(\mathcal{I}+\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}})=\mathrm{ran}\mathcal{Q}$.
The proof is exactly the same as [5, Theorem 21.1], if the subspace $\mathrm{ran}\mathcal{Q}$ is regarded as the ambient space. Actually, our specific $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}$ in Theorem 3.4 can be replaced by any monotone operator $\mathcal{B}:\mathrm{ran}\mathcal{Q}\mapsto 2^{\mathrm{ran}\mathcal{Q}}$, of course, including $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathrm{ran}\mathcal{Q}}$.